### 07 Apr 2026 v1

@@ 1 Introduction
Merger timescale prescriptions play a central role in semi-analytic models [14, 15, 16, 17], where they are used to translate the merger history of subhalos into predictions for galaxy–galaxy mergers. This connection enables SAMs to study a range of merger-driven physical processes, such as morphological transformations [18, 19, 20], starbursts [21, 22, 23, 24], and AGN activity [25, 26].
@@
Another complication stems from the hierarchical nature of accretion itself. It is common for subhalos to be accreted as members of groups (“group infall”), which naturally establishes a nested, multi-level hierarchy within the host halo [37, 38, 39, 40, 41, 42]. This hierarchy entails that merging with the central subhalo is not the only fate for accreted subhalos, and some will merge with other satellite subhalos during their orbital evolution. Although central–satellite and satellite–satellite mergers are often analyzed separately, detailed studies that track the fate of accreted groups of subhalos and quantify the satellite–satellite merger rate remain limited.
By leveraging a dark-matter-only (DMO) cosmological simulation analyzed with the hbt+ subhalo finder [43, 44], we investigate subhalo mergers driven by dynamical friction. Unlike methods that rely on thresholds measured on the scale of the subhalo (e.g., bound particle number or orbital angular momentum), hbt+ determines the phase-space coalescence of a subhalo with the dense core of another subhalo. We refer to mergers identified through this physically motivated criterion as subhalo sinking (see Section 2.2 for details). The term “sinking” is used interchangeably with “merger” throughout this work, but specifically denotes those merger events identified by the hbt+ algorithm. Established terms such as “central–satellite merger” and “major merger” retain their conventional meanings. This work aims to:
@@
Alhough the mergers among subhalos are rooted in the largely self-similar merger tree of the host halo [40], we will show that subhalo mergers are much more complex and deviates from self-similarity, due to additional physical processes during subhalo evolution since their accretion.
@@ 3 Hierarchical Sinking Pathways of Subhalos
The prevailing framework in the literature classifies subhalo mergers into two broad channels: central–satellite and satellite–satellite mergers. While conceptually clear, this binary scheme does not fully capture the complexity inherent in hierarchically structured systems. In the subhalo mergers defined via sinking, both channels encompass a diversity of sinking pathways, as satellites themselves often host their own substructures, creating a nested hierarchy.
@@
Figure 3 shows the distribution of parent levels for sunken child subhalos binned by their own level $\ell_{\rm sub}$. Our analysis reveals a clear hierarchical sinking pattern. The vast majority of subhalo sinking events occur between a subhalo and its direct parent (level $\ell_{\rm sub}-1$). This is consistent with the findings in R. E. Angulo et al. [29], where they demonstrated that subhalos involved in satellite–satellite mergers are dynamically associated with each other prior to infall, rather than forming through random encounters within the dark matter halo.
@@ 4 The Subhalo Sinking Rate
The merger rates between subhalos and galaxies have been extensively studied in the literature [34, 31, 30, 27, 32, 28, 33]. In this section, we examine the subhalo merger rate identified via subhalo sinking and revisit subhalo sinking by introducing a finer classification based on the hierarchical level $\ell$ of satellites within the subhalo structure. We focus exclusively on sinking events between child subhalos and their direct parent subhalos, as these account for the vast majority of sinking events ($90\%$). We compare the sinking rates across different groups and halos and find that satellite–satellite sinking events constitute a non-negligible fraction of the overall sunken population.
@@ 4.2 Key Results and Interpretation
Tracing the redshift evolution of $f_{\rm halo}$ above fixed mass-ratio thresholds (Figure 8), we find that $f_{\rm halo}$ increases steadily in the form of $\propto(1+z)^{\alpha}$ for the first three levels. The slope $\alpha$ varies across the hierarchical levels with $\alpha=0.13,0.61,0.82$ for $\ell=1,2,3$ separately. Notably, for level‑2 subhalos with $\mu_{\rm peak}>10^{-2}$, their sinking rate becomes comparable to that of central–satellite (level‑1) sinking events, highlighting the significant aggregate contribution of satellite–satellite mergers. It is expected that the $f_{\rm halo}$ with $\ell=3$ can also rival the central-satellite rate when less massive subhalos are resolved in the simulations with higher particle resolution. The steady rise of the halo-level sinking rate for deep-level subhalos toward lower mass thresholds reveals that satellite–satellite mergers become increasingly significant as we resolve subhalos with progressively smaller peak mass ratios.
@@ 4.3 Comparison to Previous Work
A. R. Wetzel et al. [28] report that satellite–satellite mergers constitute only a negligible fraction of the overall subhalo merger population compared with the central-satellite mergers. In contrast, we find a comparable satellite–satellite sinking rate in both Figures 5 and 8. This apparent discrepancy, however, stems from differences in definitions and sample selection, rather than indicating a fundamental inconsistency.
A. R. Wetzel et al. [28] quantify the contribution of satellite–satellite mergers by measuring their fraction among all merger events whose descendant subhalos lie within a fixed bound mass bin (e.g., $10^{12},10^{13}]\,\mathrm{M}_{\odot}\,h^{-1}$). They find that satellite–satellite mergers account for only $5\%$$15\%$ of the total. This result is a natural consequence of their sample selection: within the descendant mass bins they consider, central subhalos dominate the population, while satellites constitute only a small fraction. Consequently, most merger events involving these descendants are central–satellite mergers, and satellite–satellite mergers are inevitably subdominant. While this statistic usefully characterizes the relative abundance of satellite–satellite mergers in a descendant-selected sample, it does not directly measure how frequently subhalos of shallower levels receive merged subhalos from deeper levels, nor does it reveal how the merger rate varies across hierarchical levels.
@@
Together, these two complementary definitions demonstrate that satellite–satellite sinking events are not a negligible minority, but rather an important process that must be accounted for in any realistic model of subhalo evolution.
@@ 5 The radial distribution of the subhalo sinking events
Central–satellite sinking events, in which a subhalo sinks directly into the central (level-0) subhalo, exhibit a centrally concentrated radial profile regardless of the levels of subhalos. Whether the sinking involves a level-1, level-2, or level-3 subhalo sinking into the central, the distribution peaks at $r\leq 0.1R_{\rm vir}$ with an extended tail. The shape of this central concentration is remarkably consistent across levels, differing only in amplitude (see Figure 9). This universality reflects the common physical driver: dynamical friction operating deep within the host’s potential well. Notably, a small fraction of these central-satellite events also occur at large radii ($r\gtrsim 0.5R_{\rm vir}$). These distant events can be attributed either to central subhalos with unusually diffuse spatial distributions or, more commonly, to satellite subhalos whose assigned positions are associated with large core spatial dispersions ($\sigma_{R}$) via the reverse-link criterion. In both cases, the sunken subhalos have $\sigma_{R}$ values in the top 5% of the overall spatial dispersion distribution (Figure 11), confirming their anomalously diffuse nature.
@@
The dominance of satellite–satellite sinking events in the outer regions is consistent with previous findings [29, 64]. This suggests that subhalo groups near the center cannot retain their member subhalos long enough for orbital decay to complete.
@@
Taken together, the radial probabilities highlight that the subhalo sinking is intimately connected to the hierarchical nature of structure formation. While central–satellite mergers dominate the innermost halo regions, satellite–satellite mergers preferentially occupy the outskirts, underscoring the key roles played by group infall, tidal perturbations, and the multi-level subhalo hierarchy in governing the locations of subhalo coalescence.
@@ 6.1 Resolution Dependence
Figure 11: Upper panels: Spatial dispersion of the 20 most bound particles of subhalos at $z=0$ as a function of bound mass. The left and right panels correspond to central and satellite subhalos, respectively. The blue solid line indicates the median dispersion in each mass bin, and the black dashed line marks the simulation softening length. The region between the two gray dashed lines in each panel encloses the central 90% (5th–95th percentile) of subhalos at a given mass. For clarity, only $10^{5}$ randomly chosen subhalo samples are shown, with colors indicating subhalo density. Lower panels: Velocity dispersion of the 20 most bound particles of subhalos, following the same layout as the upper panels. The blue solid line represents the median of the overall velocity dispersion distribution, and the black dashed line shows the $\sigma_{v}-N_{\rm particle}$ fitting formula derived from the Millennium-II simulation in [44]. As above, the gray dashed lines enclose the 5th–95th percentile range in each mass bin.
@@ 6.2.1 Classification of Sinking Population
  • Region 2 (child‑dispersion–driven sinking): Here, $\min\delta_{\rm s,c}<2\leq\min\delta_{\rm s,p}$. This population is uniquely recovered by the refined algorithm and accounts for about 24.2% of the total sunken population. It is mainly contributed by child subhalos that exhibit larger internal spatial dispersion ($\sigma_{x,c}$) relative to their parents (see Figure 2).
  • Region 3 (parent‑dispersion–driven sinking): A minority of events (16.0%) satisfy $\min\delta_{\rm s,p}<2$ while $\min\delta_{\rm s,c}\geq 2$, meaning they are identified solely by the $\delta_{\rm s,p}$ criterion. This typically occurs when the parent subhalo has a slightly larger internal velocity or spatial dispersion than its child, causing $\delta_{\rm s,p}$ to become the smaller distance and fall just below the threshold. Representative evolutionary tracks for this sinking type are shown in Figure 15.
Overall, the refined algorithm recovers approximately 32% more sinking events than the original single‑direction version, with the addition dominated by child‑dispersion–driven sinking.
@@ 6.2.2 Driving Mechanisms: Orbital Decay versus Tidal Heating
We have shown representative case studies for Regions 1, 2, and 3 in Figures 12, and 15, respectively. Specifically, in reverse‑link cases, tidal heating appears to be the dominant mechanism driving the sinking process, causing $\delta_{\rm s,c}$ to drop below the threshold. In contrast, parent‑dispersion–driven and doubly identified sunken subhalos show less evidence of significant heating; their cores remain compact throughout the sinking process.
@@
In child‑dispersion–driven cases, the majority ($70\%$) of sunken satellite subhalos experience significant tidal heating, with their $\sigma^{2}_{R}$ values increasing by a factor of $\sim 2$ after sinking, while their parent subhalos remain nearly intact and compact. This confirms that reverse‑link sinking events are primarily driven by heating at the interface between the two evolutionary phases. A direct consequence of the substantial tidal heating is that the child subhalo quickly disrupts in about 1-2 dynamical times, which happens to the majority of the child-dispersion-driven cases.
Conversely, in parent‑dispersion–driven and doubly identified sinking events, the sunken satellites remain compact throughout. Although the spatial dispersion of their parent subhalos may expand slightly, the child subhalos themselves show no evidence of tidal heating. In these cases, the transition between the two‑stage evolution is driven primarily by orbital decay rather than by heating.
@@ 7 Summary and Conclusions
    Satellite–satellite mergers preferentially occur in the halo outskirts. In contrast to central–satellite mergers, which occur near the halo center, Figure 9 shows that sinking events of the deeper-level subhalos preferentially occur at larger radii.

### 08 Apr 2026 v2

@@ 1 Introduction
Merger timescale prescriptions play a central role in semi-analytic models [14, 15, 16, 17], where they are used to translate the merger history of subhalos into predictions for galaxy-galaxy mergers. This connection enables SAMs to study a range of merger-driven physical processes, such as morphological transformations [18, 19, 20], starbursts [21, 22, 23, 24], and AGN activity [25, 26].
@@
Another complication stems from the hierarchical nature of accretion itself. It is common for subhalos to be accreted as members of groups (“group infall”), which naturally establishes a nested, multi-level hierarchy within the host halo [37, 38, 39, 40, 41, 42]. This hierarchy entails that merging with the central subhalo is not the only fate for accreted subhalos, and some will merge with other satellite subhalos during their orbital evolution. Although central-satellite and satellite-satellite mergers are often analyzed separately, detailed studies that track the fate of accreted groups of subhalos and quantify the satellite-satellite merger rate remain limited.
By leveraging a dark-matter-only (DMO) cosmological simulation analyzed with the hbt+ subhalo finder [43, 44], we investigate subhalo mergers driven by dynamical friction. Unlike methods that rely on thresholds measured on the scale of the subhalo (e.g., bound particle number or orbital angular momentum), hbt+ determines the phase-space coalescence of a subhalo with the dense core of another subhalo. We refer to mergers identified through this physically motivated criterion as subhalo sinking (see Section 2.2 for details). The term “sinking” is used interchangeably with “merger” throughout this work, but specifically denotes those merger events identified by the hbt+ algorithm. Established terms such as “central-satellite merger” and “major merger” retain their conventional meanings. This work aims to:
@@
Although the mergers among subhalos are rooted in the largely self-similar merger tree of the host halo [40], we will show that subhalo mergers are much more complex and deviate from self-similarity, due to additional physical processes during subhalo evolution since their accretion.
@@ 3 Hierarchical Sinking Pathways of Subhalos
The prevailing framework in the literature classifies subhalo mergers into two broad channels: central-satellite and satellite-satellite mergers. While conceptually clear, this binary scheme does not fully capture the complexity inherent in hierarchically structured systems. In the subhalo mergers defined via sinking, both channels encompass a diversity of sinking pathways, as satellites themselves often host their own substructures, creating a nested hierarchy.
@@
Figure 3 shows the distribution of parent levels for sunken child subhalos binned by their own level $\ell_{\rm sub}$. Our analysis reveals a clear hierarchical sinking pattern. The vast majority of subhalo sinking events occur between a subhalo and its direct parent (level $\ell_{\rm sub}-1$). This is consistent with the findings in R. E. Angulo et al. [29], where they demonstrated that subhalos involved in satellite-satellite mergers are dynamically associated with each other prior to infall, rather than forming through random encounters within the dark matter halo.
@@ 4 The Subhalo Sinking Rate
The merger rates between subhalos and galaxies have been extensively studied in the literature [34, 31, 30, 27, 32, 28, 33]. In this section, we examine the subhalo merger rate identified via subhalo sinking and revisit subhalo sinking by introducing a finer classification based on the hierarchical level $\ell$ of satellites within the subhalo structure. We focus exclusively on sinking events between child subhalos and their direct parent subhalos, as these account for the vast majority of sinking events ($90\%$). We compare the sinking rates across different groups and halos and find that satellite-satellite sinking events constitute a non-negligible fraction of the overall sunken population.
@@ 4.2 Key Results and Interpretation
Tracing the redshift evolution of $f_{\rm halo}$ above fixed mass-ratio thresholds (Figure 8), we find that $f_{\rm halo}$ increases steadily in the form of $\propto(1+z)^{\alpha}$ for the first three levels. The slope $\alpha$ varies across the hierarchical levels with $\alpha=0.13,0.61,0.82$ for $\ell=1,2,3$ separately. Notably, for level‑2 subhalos with $\mu_{\rm peak}>10^{-2}$, their sinking rate becomes comparable to that of central-satellite (level‑1) sinking events, highlighting the significant aggregate contribution of satellite-satellite mergers. It is expected that the $f_{\rm halo}$ with $\ell=3$ can also rival the central-satellite rate when less massive subhalos are resolved in the simulations with higher particle resolution. The steady rise of the halo-level sinking rate for deep-level subhalos toward lower mass thresholds reveals that satellite-satellite mergers become increasingly significant as we resolve subhalos with progressively smaller peak mass ratios.
@@ 4.3 Comparison to Previous Work
A. R. Wetzel et al. [28] report that satellite-satellite mergers constitute only a negligible fraction of the overall subhalo merger population compared with the central-satellite mergers. In contrast, we find a comparable satellite-satellite sinking rate in both Figures 5 and 8. This apparent discrepancy, however, stems from differences in definitions and sample selection, rather than indicating a fundamental inconsistency.
A. R. Wetzel et al. [28] quantify the contribution of satellite-satellite mergers by measuring their fraction among all merger events whose descendant subhalos lie within a fixed bound mass bin (e.g., $10^{12},10^{13}]\,\mathrm{M}_{\odot}\,h^{-1}$). They find that satellite-satellite mergers account for only $5\%$-$15\%$ of the total. This result is a natural consequence of their sample selection: within the descendant mass bins they consider, central subhalos dominate the population, while satellites constitute only a small fraction. Consequently, most merger events involving these descendants are central-satellite mergers, and satellite-satellite mergers are inevitably subdominant. While this statistic usefully characterizes the relative abundance of satellite-satellite mergers in a descendant-selected sample, it does not directly measure how frequently subhalos of shallower levels receive merged subhalos from deeper levels, nor does it reveal how the merger rate varies across hierarchical levels.
@@
Together, these two complementary definitions demonstrate that satellite-satellite sinking events are not a negligible minority, but rather an important process that must be accounted for in any realistic model of subhalo evolution.
@@ 5 The radial distribution of the subhalo sinking events
Central-satellite sinking events, in which a subhalo sinks directly into the central (level-0) subhalo, exhibit a centrally concentrated radial profile regardless of the levels of subhalos. Whether the sinking involves a level-1, level-2, or level-3 subhalo sinking into the central, the distribution peaks at $r\leq 0.1R_{\rm vir}$ with an extended tail. The shape of this central concentration is remarkably consistent across levels, differing only in amplitude (see Figure 9). This universality reflects the common physical driver: dynamical friction operating deep within the host’s potential well. Notably, a small fraction of these central-satellite events also occur at large radii ($r\gtrsim 0.5R_{\rm vir}$). These distant events can be attributed either to central subhalos with unusually diffuse spatial distributions or, more commonly, to satellite subhalos whose assigned positions are associated with large core spatial dispersions ($\sigma_{R}$) via the reverse-link criterion. In both cases, the sunken subhalos have $\sigma_{R}$ values in the top 5% of the overall spatial dispersion distribution (Figure 11), confirming their anomalously diffuse nature.
@@
The dominance of satellite-satellite sinking events in the outer regions is consistent with previous findings [29, 64]. This suggests that subhalo groups near the center cannot retain their member subhalos long enough for orbital decay to complete.
@@
Taken together, the radial probabilities highlight that the subhalo sinking is intimately connected to the hierarchical nature of structure formation. While central-satellite mergers dominate the innermost halo regions, satellite-satellite mergers preferentially occupy the outskirts, underscoring the key roles played by group infall, tidal perturbations, and the multi-level subhalo hierarchy in governing the locations of subhalo coalescence.
@@ 6.1 Resolution Dependence
Figure 11: Upper panels: Spatial dispersion of the 20 most bound particles of subhalos at $z=0$ as a function of bound mass. The left and right panels correspond to central and satellite subhalos, respectively. The blue solid line indicates the median dispersion in each mass bin, and the black dashed line marks the simulation softening length. The region between the two gray dashed lines in each panel encloses the central 90% (5th-95th percentile) of subhalos at a given mass. For clarity, only $10^{5}$ randomly chosen subhalo samples are shown, with colors indicating subhalo density. Lower panels: Velocity dispersion of the 20 most bound particles of subhalos, following the same layout as the upper panels. The blue solid line represents the median of the overall velocity dispersion distribution, and the black dashed line shows the $\sigma_{v}-N_{\rm particle}$ fitting formula derived from the Millennium-II simulation in [44]. As above, the gray dashed lines enclose the 5th-95th percentile range in each mass bin.
@@ 6.2.1 Classification of Sinking Population
  • Region 2 (child‑dispersion-driven sinking): Here, $\min\delta_{\rm s,c}<2\leq\min\delta_{\rm s,p}$. This population is uniquely recovered by the refined algorithm and accounts for about 24.2% of the total sunken population. It is mainly contributed by child subhalos that exhibit larger internal spatial dispersion ($\sigma_{x,c}$) relative to their parents (see Figure 2).
  • Region 3 (parent‑dispersion-driven sinking): A minority of events (16.0%) satisfy $\min\delta_{\rm s,p}<2$ while $\min\delta_{\rm s,c}\geq 2$, meaning they are identified solely by the $\delta_{\rm s,p}$ criterion. This typically occurs when the parent subhalo has a slightly larger internal velocity or spatial dispersion than its child, causing $\delta_{\rm s,p}$ to become the smaller distance and fall just below the threshold. Representative evolutionary tracks for this sinking type are shown in Figure 15.
Overall, the refined algorithm recovers approximately 32% more sinking events than the original single‑direction version, with the addition dominated by child‑dispersion-driven sinking.
@@ 6.2.2 Driving Mechanisms: Orbital Decay versus Tidal Heating
We have shown representative case studies for Regions 1, 2, and 3 in Figures 12, and 15, respectively. Specifically, in reverse‑link cases, tidal heating appears to be the dominant mechanism driving the sinking process, causing $\delta_{\rm s,c}$ to drop below the threshold. In contrast, parent‑dispersion-driven and doubly identified sunken subhalos show less evidence of significant heating; their cores remain compact throughout the sinking process.
@@
In child‑dispersion-driven cases, the majority ($70\%$) of sunken satellite subhalos experience significant tidal heating, with their $\sigma^{2}_{R}$ values increasing by a factor of $\sim 2$ after sinking, while their parent subhalos remain nearly intact and compact. This confirms that reverse‑link sinking events are primarily driven by heating at the interface between the two evolutionary phases. A direct consequence of the substantial tidal heating is that the child subhalo quickly disrupts in about 1-2 dynamical times, which happens to the majority of the child-dispersion-driven cases.
Conversely, in parent‑dispersion-driven and doubly identified sinking events, the sunken satellites remain compact throughout. Although the spatial dispersion of their parent subhalos may expand slightly, the child subhalos themselves show no evidence of tidal heating. In these cases, the transition between the two‑stage evolution is driven primarily by orbital decay rather than by heating.
@@ 7 Summary and Conclusions
    Satellite-satellite mergers preferentially occur in the halo outskirts. In contrast to central-satellite mergers, which occur near the halo center, Figure 9 shows that sinking events of the deeper-level subhalos preferentially occur at larger radii.